OpenStudy (angelwings996):

For the function f(x) = (3 - 4x)^2, find f^(-1). Determine whether f^(-1) is a function.

OpenStudy (angelwings996):

@wio

OpenStudy (anonymous):

Solve for \(x\).

OpenStudy (angelwings996):

Wouldn't I plug -1 in for x ?

OpenStudy (anonymous):

Nope. They want you to find the inverse.

OpenStudy (anonymous):

no no no. That you would do when you are asked for f(-1) But there you are asked for the inverse function. so u solve for x

OpenStudy (angelwings996):

Okay, so... would it be (x - 3/4)^2 ?

OpenStudy (anonymous):

y = (3-4x)^2 sqrt(y) = 3-4x 4x = 3 - sqrt(y) x = (1/4)(3 - sqrt(y)) Therefore f^-1(x) is \[(3-\sqrt{x}) / 4\]

OpenStudy (angelwings996):

Okay, so would that make f^(-1) a function? @Saikam

OpenStudy (anonymous):

no. because it gives you 2 different values for single x. Technically, its not a function.

jimthompson5910 (jim_thompson5910):

If y^2 = x then y = sqrt(x) or y = -sqrt(x) This means that y^2 = x is not a function because plugging in positive x values (say x = 4) gives you more than one y value.

OpenStudy (angelwings996):

Okay, last question for tonight I promise! What value completes the square for the expression? \[x ^{2} - 12x\]

OpenStudy (angelwings996):

I know that x^2 + 12 is 36, would this be -36 or would it change the whole thing since this is subtraction?

OpenStudy (angelwings996):

I meant x^2 + 12x is 36, sorry

jimthompson5910 (jim_thompson5910):

Similar Problem: x^2 - 8x take half of -8 to get -4, square -4 to get 16 so you add 16 to the expression above to make it a perfect square so x^2 - 8x + 16 is a perfect square

OpenStudy (angelwings996):

well it is a multiple chice question. It's either 6, -6, 36 or -36

OpenStudy (angelwings996):

*choice

jimthompson5910 (jim_thompson5910):

do you see how I got 16 in the example problem x^2 - 8x ?

OpenStudy (angelwings996):

from 4^2 ?

jimthompson5910 (jim_thompson5910):

(-4)^2 actually, you're close

OpenStudy (angelwings996):

Oh okay

jimthompson5910 (jim_thompson5910):

you follow the same steps to find the answer to this problem

OpenStudy (angelwings996):

so half of -12 is -6 right? so it would be -6^2 which would be 36?

jimthompson5910 (jim_thompson5910):

(-6)^2 = 36, yes you got it right

OpenStudy (angelwings996):

Okay, thank you so much for all of your help! Have a good night everyone (: